i need help with the following questions:

1) -17 = 19 + x


2) g - 35 = -62


3) -27 + z = 47

Answer by rmromero(383)

Solution: We are going to solve for the variable asked.
1) -17 = 19 + x
To solve for x, we must separate the variable and the constant. We must eliminate 19 so that x will be left on the right side. 19 is positive so we will ADD the Negative Inverse of 19 both sides. That will give us -19.


-17 + (-19) = 19 + x + (-19)
-17 - 19 = 19 + x - 19

-36 = x



2) g - 35 = -62
Lets do the same, Add Negative Inverse of -35 both sides. That is 35


g - 35 + 35 = -62 + 35

g = -27



3) -27 + z = 47
Add negative inverse of -27 both sides, that is 27


-27 + z + 27 = 47 + 27

z = 74